OpenStudy (anonymous):

Find the solution of the IVP 2x^2y"+xy'-3y=0 y(1)=1, y'(1)=4 Describe how the solution behaves as x goes to zero.

OpenStudy (anonymous):

How far did you get on this one?

OpenStudy (anonymous):

i assumed that the power series would be centred on 1, so i let t=x-1 and z(t)=y(x) then i found the power series expansion, and found that a2=a0/2 a3=a0/3 a4=-11a0/24 a5=-41a0/60 but i don't think thats right…. I am trying to relate the expressions so I can write a general solution for y(x0=… but I don't see how the a coefficient expressions relate.

OpenStudy (anonymous):

Let y=x^m

OpenStudy (anonymous):

y'=mx^(m-1)

OpenStudy (anonymous):

y''=m(m-1)x^(m-2)

OpenStudy (anonymous):

Substitute these into the DE and solve for m.

OpenStudy (anonymous):

Don't go by the power series method. What do you get for m?

OpenStudy (anonymous):

Reason why you do not need to go by the series method for this DE is because it is a Cauchy Euler DE. For Cauchy Euler DE, letting y=x^m is a simpler approach.

OpenStudy (anonymous):

we have to use power series to solve this because that is the section we are learning in class!

OpenStudy (anonymous):

in the question it says we have to use power series

OpenStudy (anonymous):

OK. Then I would just let \[y=\sum_{n=0}^{\infty}a _{n}x ^{n}\]

OpenStudy (anonymous):

The final answer to the question as to how it should behave as x goes to 0, is y goes to e.

OpenStudy (anonymous):

could you maybe write out the first few steps? I don't get what you mean when you say y goes to e is the answer

OpenStudy (anonymous):

Let me work on this. Will revert.

OpenStudy (anonymous):

Here is the solution.

OpenStudy (anonymous):

The corret response as x tends to 0, is y tends to - infinity.

OpenStudy (anonymous):

I took a look at your solution, is there any way you can center around 1 and solve it that way?

OpenStudy (anonymous):

Sorry I dont know why you want to do that and I am not about to d that. Cauchy Euler DE is a very standard one.